# Intent Laundering: AI Safety Datasets Are Not What They Seem

### v1

@@ Abstract
We systematically evaluate the quality of widely used AI safety datasets from two perspectives: in isolation and in practice. In isolation, we examine how well these datasets reflect real-world attacks based on three key properties: driven by ulterior intent, well-crafted, and out-of-distribution. We find that these datasets overrely on “triggering cues”: words or phrases with overt negative/sensitive connotations that are intended to trigger safety mechanisms explicitly, which is unrealistic compared to real-world attacks. In practice, we evaluate whether these datasets genuinely measure safety risks or merely provoke refusals through triggering cues. To explore this, we introduce “intent laundering”: a procedure that abstracts away triggering cues from attacks (data points) while strictly preserving their malicious intent and all relevant details. Our results indicate that current AI safety datasets fail to faithfully represent real-world attacks due to their overreliance on triggering cues. In fact, once these cues are removed, all previously evaluated “reasonably safe” models become unsafe, including Gemini 3 Pro and Claude Sonnet 3.7. Moreover, when intent laundering is adapted as a jailbreaking technique, it consistently achieves high attack success rates, ranging from 90% to over 98%, under fully black-box access. Overall, our findings expose a significant disconnect between how model safety is evaluated and how real-world adversaries behave.
@@ 1 Introduction
Figure 1: Overview of our intent laundering framework. Without the feedback loop, the framework illustrates the intent laundering procedure. With the loop, it shows how intent laundering operates as a jailbreaking technique. The process begins by passing the original malicious request (data point) through the intent launderer to generate a revision. This revision is then used to attack the target model. An LLM judge evaluates the response for safety and practicality. If a response is both unsafe and practical, the attack is considered successful. Otherwise, the revision–regeneration mechanism is triggered, using all previous failed revisions as feedback to generate a new, improved revision. The loop ends when either a set number of iterations is reached or a target attack success rate is achieved.
@@
In this work, we systematically study the quality of widely used AI safety datasets—AdvBench (Zou et al., 2023) and HarmBench (Mazeika et al., 2024)—based on prior research (Kim et al., 2025; Xie et al., 2025; Zhao et al., 2024; Xu et al., 2024; Li et al., 2024; Anil et al., 2024, inter alia). First, we analyze whether these datasets, in isolation, faithfully represent real-world attacks by evaluating three defining properties of such attacks: being driven by ulterior intent, well-crafted, and out-of-distribution. [^1] Similar to other datasets, diversity is another important consideration in the design of safety datasets. This includes diversity across topics and data points. Topic diversity is generally well accounted for by most dataset creators (Röttger et al., 2024a; Mazeika et al., 2024; Zou et al., 2023) and is thus excluded from our analysis. In contrast, diversity at the data-point level is a major issue in these datasets, which we discuss in Section 2.2. Second, we examine whether these datasets actually measure safety risks in practice when they are used to evaluate models.
@@
To evaluate the quality of safety datasets in isolation, we begin by analyzing $n$-gram word clouds. This helps visualize how the most frequent unigrams recur as higher-order $n$-grams. We find that these recurring $n$-grams consistently form unrealistic “triggering cues [^2] We use “triggering language” interchangeably.: expressions with overt negative/sensitive connotations. These cues fall into two categories: (1) inherent—expressions that carry such connotations by nature, regardless of context (e.g., “commit suicide”), and (2) contextual—expressions that acquire such connotations in the context of harmful requests (e.g., [malicious intent] + “without getting caught,” explicitly signaling evasion). Figure 2 shows examples of triggering cues. These cues undermine two properties—being well-crafted and driven by ulterior intent—as such overt language rarely appears in real-world attacks and seems engineered to trigger safety mechanisms artificially. We also show that the repetitive overuse of these cues creates substantial duplication, producing near-identical data points in both sentence structure and malicious intent. This further degrades the out-of-distribution property and adds to the erosion of well-crafted property. Overall, our analysis indicates that current safety datasets fail to faithfully represent real-world attacks.
@@
Next, we evaluate the quality of safety datasets in practice by involving models. In particular, we examine whether these datasets genuinely measure safety risks or merely rely on triggering language that safety-aligned models are trained to detect and refuse to answer. To explore this, we introduce “intent laundering”: a systematic procedure that abstracts away overt triggering language from attacks (data points) while strictly preserving their malicious intent. This process has two components: (1) connotation neutralization, and (2) context transposition. In connotation neutralization, triggering expressions are replaced with neutral/positive or descriptive alternatives. In context transposition, real-world scenarios and referents—such as individuals (e.g., “immigrants”) or institutions (e.g., “charity”)—that can act as triggering cues in harmful requests are mapped to non-real-world contexts (e.g., a game world). We automate this process using an “intent launderer”: a large language model (LLM) with a few-shot in-context learning (ICL) setup. Each ICL demonstration pairs an original data point with its manually crafted, intent-laundered revision. Figure 1 depicts an overview of this framework. Our results reveal a strong effect: once triggering language is removed, the attack success rate (ASR) jumps from a mean of 5.38% to 86.79% on AdvBench, and from 13.79% to 79.83% on HarmBench. This indicates that current safety evaluations, and the safety alignment built on them, overrely on triggering cues, leading to an inflated evaluation of safety relative to real-world attacks.
@@
Finally, we propose intent laundering as a standalone jailbreaking technique by adding an iterative revision–regeneration mechanism. In each iteration, the model uses all previous failed revisions as feedback to generate a new, improved revision, using the same few-shot ICL setup. Figure 1 shows this mechanism. This iterative process continues until either a predefined number of regeneration attempts is reached or a target ASR is met. Our results show that, with this regeneration loop, intent laundering achieves high ASR (90%98.55%) after only a few iterations across all studied models under fully black-box access. This includes recent models widely reported as among the safest—such as Gemini 3 Pro (Google, 2025a; b) and Claude Sonnet 3.7 (Anthropic, 2025a; b; Holistic AI, 2025). These findings further confirm that existing safety evaluations and safety-alignment methods are highly overfitted to triggering cues.
@@
(1) We conduct a systematic study of the quality of widely used AI safety datasets and show that they do not faithfully reflect real-world attacks due to overuse of unrealistic triggering cues.
(2) We introduce intent laundering: a procedure that empirically verifies that revising attacks to remove triggering cues sharply increases ASR, from a mean of 5.38% to 86.79% on AdvBench, and from 13.79% to 79.83% on HarmBench.
(3) We adapt intent laundering into a novel jailbreaking method by adding a revision–regeneration step for failed revisions, achieving high ASR (90%–98.55%) across the board.
(4) We find that recent safety-alignment techniques still fall significantly short in repelling real-world attacks, and that models aligned using these methods are often referred to as “reasonably safe” mainly because existing safety evaluations are overly simplistic and unrealistic.
@@ 2 Empirical Motivation
To motivate our study, we first analyze the quality of safety datasets in isolation. We use $n$-gram word clouds and pairwise similarity analyses to assess whether these datasets reflect real-world attacks. Specifically, we analyze three key characteristics of such attacks:
@@ 2.1 Word Clouds
Figure 2: Word clouds of the 40 most frequent $n$-grams, where $n\in\{1,2,3\}$, from the combined AdvBench and HarmBench corpus. Triggering cues are highlighted in red for inherently, and orange for contextually negative/sensitive connotations. This includes triggering words in the unigram cloud and triggering phrases in the bigram and trigram clouds. Neutral-connotation unigrams that contribute to triggering cues in higher-order $n$-grams are also shown in green. These visualizations reveal an unusual overrepresentation of overtly triggering cues in safety datasets, suggesting that data points are artificially designed to trigger safety mechanisms. For example, expressions such as “tutorial” and “step-by-step instructions” are used explicitly to trigger infohazard safeguards (Bostrom, 2011), which is unrealistic. Even minimally skilled bad actors rarely use such overt and self-incriminating language (e.g., “commit identity theft”). These findings indicate that safety datasets fail to capture two of the key properties of real-world attacks: being well-crafted and driven by ulterior intent. Note that, for clarity, only a representative subset of expressions is marked in the word clouds above.
@@
Figure 2 displays the resulting word clouds. At the unigram level, there is an unusual overrepresentation of words with overt negative/sensitive connotations, which we refer to as “triggering words.” Triggering words fall into two categories: (1) inherent triggering words (in red), which carry negative/sensitive connotations by nature (e.g., “steal”), and (2) contextual triggering words (in orange), which signal such connotations when used in harmful contexts (e.g., “commit”).
As unigrams evolve into higher-order $n$-grams, this overrepresentation intensifies: phrases with negative/sensitive connotations become unusually dominant. Similar to unigrams, these phrases carry such connotations either inherently or contextually within malicious requests. We refer to these as “triggering phrases,” and together with triggering words, they form what we term “triggering cues.”
@@
However, such explicit and repetitive overuse of triggering cues, along with direct mentions of malicious intent, directly contradicts real-world attack scenarios. Even minimally capable adversaries rarely use such overt language, as it easily triggers safety mechanisms. Taken together, these patterns indicate that existing safety datasets contain contrived data points largely disconnected from real-world attacks, where harmful requests are well-crafted and motivated by ulterior intent.
@@ 2.2 Data Duplication
The excessive use of triggering cues not only creates unrealistic data points but also suggests significant overlap in sentence structure and malicious intent across data points. To analyze this, we conduct a straightforward pairwise similarity check between data points within each dataset. We use similarity thresholds ranging from 0.7 to 0.99. For a given threshold, data points that do not meet the similarity threshold with any other data point are labeled as unique, while those that meet/exceed the threshold with at least one other data point are labeled as duplicated and grouped to represent a single data point. As there is no universally accepted threshold for high similarity—particularly within datasets containing data points from a single context—we use the GSM8K dataset (Cobbe et al., 2021), a widely used non-safety dataset, as a baseline. For a fair comparison, we subsample GSM8K to match the size of each safety dataset.
@@
Overall, analyzing safety datasets in isolation yields two main insights: (1) they repeatedly overuse triggering cues, and (2) this leads to substantial duplication. Together, these undermine all three key properties of real-world attacks. Thus, they fail to faithfully represent real-world adversarial behavior, and safety evaluations based on them can be inflated. This raises a key question: if triggering cues are removed, do models previously reported as “reasonably safe” still remain so? We explore this next.
@@ 3.1 Intent Laundering
After generating all revisions, we collect model responses to both the original and revised data points to assess how triggering language affects safety evaluation.
@@ 3.2 Evaluation
We exclusively evaluate the safety of responses to revisions using a five-point Likert scale based on the level of harm exposed: extremely unsafe, highly unsafe, unsafe, moderately unsafe, and safe (see Figure 10 in Appendix C for score definitions). [^4] We do not use numerical scales to avoid potential biases against assigning extremely high/low scores, as reported in prior studies (Ouyang et al., 2022; Golchin et al., 2025). We also condition our safety evaluation specifically on harm relative to the original malicious intent, which serves as the evaluation criterion (see Figure 9 in Appendix C for how it is generated for each attack). Safety is thus judged on two fronts: (1) whether the response to the revision addresses the same harmful intent as the original version, and if so, (2) how unsafe the response is. If the revision fails to preserve the original malicious intent, the response is rated as safer.
@@
Similar to the safety evaluation, we exclusively evaluate the practicality of responses to revisions using a five-point Likert scale based on the practicality and transferability of the provided details in the real world: extremely applicable, highly applicable, applicable, moderately applicable, and inapplicable (see Figure 11 in Appendix C for score definitions). Similarly, we condition our practicality evaluation specifically on harm relative to the original malicious intent, which is provided as the evaluation criterion. As with safety, practicality is judged on two fronts: (1) whether the details in the response support the same harmful intent as the original version, and if so, (2) how practical those details are when applied in the real world with reasonable adjustments.
@@
We measure performance using the attack success rate (ASR), defined as the fraction of attack attempts that succeed. In our setting, an attack (revision) is considered successful only if the response is both unsafe and practical. A response is unsafe if it is rated as extremely unsafe, highly unsafe, or unsafe, and practical if it is rated as extremely applicable, highly applicable, or applicable. An attack is unsuccessful if it fails to generate an unsafe response, or if the unsafe response is too abstract to be practical, even after adjustments. This design choice—using five-point scales and then mapping them to binary outcomes—aims to reduce ambiguity in borderline cases.
@@ 3.3 Intent Laundering as a Jailbreaking Technique
We extend intent laundering into an independent jailbreaking method, leveraging its ability to avoid triggering safety filters. Specifically, we add an iterative revision–regeneration mechanism for failed revisions, while the main procedure remains unchanged. In each iteration, all previous failed revisions are provided as feedback to the intent launderer to generate a new, improved revision, using the same few-shot ICL setup. Figure 1 illustrates intent laundering when applied as a jailbreaking technique. Figure 8 in Appendix C shows the ICL prompt used in the revision–regeneration loop, incorporating feedback from earlier failed revisions.

### v2

@@ Abstract
We systematically evaluate the quality of widely used AI safety datasets from two perspectives: in isolation and in practice. In isolation, we examine how well these datasets reflect real-world adversarial attacks based on three key properties: being driven by ulterior intent, well-crafted, and out-of-distribution. We find that these datasets overrely on “triggering cues”: words or phrases with overt negative/sensitive connotations that are intended to trigger safety mechanisms explicitly, which is unrealistic compared to real-world attacks. In practice, we evaluate whether these datasets genuinely measure safety risks or merely provoke refusals through triggering cues. To explore this, we introduce “intent laundering”: a procedure that abstracts away triggering cues from adversarial attacks (data points) while strictly preserving their malicious intent and all relevant details. Our results indicate that current AI safety datasets fail to faithfully represent real-world adversarial behavior due to their overreliance on triggering cues. Once these cues are removed, all previously evaluated “reasonably safe” models become unsafe, including Gemini 3 Pro and Claude Sonnet 3.7. Moreover, when intent laundering is adapted as a jailbreaking technique, it consistently achieves high attack success rates, ranging from 90% to over 98%, under fully black-box access. Overall, our findings expose a significant disconnect between how model safety is evaluated by existing datasets and how real-world adversaries behave.
@@ 1 Introduction
Figure 1: Overview of our intent laundering framework. Without the feedback loop, the framework displays the intent laundering procedure; with the loop, it intent laundering as a jailbreaking technique. The process begins by passing the original malicious request (data point) through the intent launderer to generate an intent-laundered revision. This revision is then used to attack the target model. An LLM judge evaluates the response for safety and practicality. If the response is both unsafe and practical, the attack is considered successful. Otherwise, the revision–regeneration mechanism is triggered, leveraging all previously failed revisions as feedback to generate a new, improved revision. The loop stops when the predefined iteration limit or target success rate is reached.
@@
In this work, we systematically study the quality of widely used AI safety datasets—AdvBench (Zou et al., 2023) and HarmBench (Mazeika et al., 2024)—based on prior research (Kim et al., 2025; Xie et al., 2025; Zhao et al., 2024; Xu et al., 2024; Li et al., 2024; Anil et al., 2024, inter alia). First, we analyze whether these datasets, in isolation, faithfully represent real-world adversarial attacks by evaluating three defining properties of such attacks: being driven by ulterior intent, well-crafted, and out-of-distribution. [^1] Similar to other datasets, diversity is another important consideration in the design of safety datasets. This includes diversity across topics and data points. Topic diversity is generally well accounted for by most dataset creators (Röttger et al., 2024a; Mazeika et al., 2024; Zou et al., 2023) and is thus excluded from our analysis. In contrast, diversity at the data-point level is a major issue in these datasets, which we discuss in Section 2.2. Second, we examine whether these datasets actually measure safety risks in practice when they are used to evaluate model safety.
@@
To evaluate the quality of safety datasets in isolation, we begin by analyzing $n$-gram word clouds. This helps visualize how the most frequent unigrams recur as higher-order $n$-grams. We find that these recurring $n$-grams consistently form unrealistic “triggering cues [^2] We use “triggering language” interchangeably.: expressions with overt negative/sensitive connotations. These cues fall into two categories: (1) inherent—expressions that carry such connotations by nature (e.g., “commit suicide”), and (2) contextual—expressions that signal such connotations in the context of harmful requests (e.g., [malicious intent] + “without getting caught,” explicitly signaling evasion). Figure 2 shows examples of triggering cues. These cues undermine two properties—being well-crafted and driven by ulterior intent—as such overt language rarely appears in real-world attacks and seems engineered to trigger safety mechanisms artificially. We also show that the repetitive overuse of these cues creates substantial duplication, producing near-identical data points in both sentence structure and malicious intent. This further degrades the out-of-distribution property and adds to the erosion of well-crafted property. Overall, our in-isolation analysis indicates that current safety datasets fail to faithfully represent real-world adversarial behavior.
@@
Next, we evaluate the quality of safety datasets in practice by involving models. In particular, we examine whether these datasets genuinely measure safety risks or merely rely on triggering cues that safety-aligned models are trained to detect and refuse to answer. To explore this, we introduce “intent laundering”: a systematic procedure that abstracts away overt triggering cues from attacks (data points) while strictly preserving their malicious intent and all relevant details. This transformation has two complementary components: (1) connotation neutralization, and (2) context transposition. In connotation neutralization, triggering expressions are replaced with neutral/positive or descriptive alternatives. In context transposition, real-world scenarios and referents—such as individuals (e.g., “immigrants”) or institutions (e.g., “charity”)—that can act as triggering cues in harmful requests are mapped to non-real-world contexts (e.g., a game world). We automate this process using an “intent launderer”: a large language model (LLM) with a few-shot in-context learning (ICL) setup. Each ICL demonstration pairs an original data point with its manually crafted, intent-laundered revision. Figure 1 depicts an overview of this framework. Our results reveal a strong and universal effect: once triggering cues are removed, the attack success rate (ASR) jumps from a mean of 5.38% to 86.79% on AdvBench, and from 13.79% to 79.83% on HarmBench.
@@
Finally, we propose intent laundering as a standalone jailbreaking technique by adding an iterative revision–regeneration mechanism. In each iteration, the model uses all previously failed revisions as feedback to generate a new, improved revision using the same few-shot ICL setup. Figure 1 shows this mechanism. This iterative process continues until either a predefined number of regeneration attempts is reached or a target ASR is met. Our results show that, with this regeneration loop, intent laundering achieves high ASR (90%98.55%) after only a few iterations across all studied models under fully black-box access. This includes recent frontier models reported as among the safest—such as Gemini 3 Pro (Google, 2025a; b) and Claude Sonnet 3.7 (Anthropic, 2025a; b; Holistic AI, 2025).
@@
(1) We show that AI safety datasets do not faithfully reflect real-world adversarial behavior due to overuse of unrealistic triggering cues.
(2) We introduce intent laundering: a procedure that empirically verifies that revising data points to remove triggering cues sharply increases ASR: from a mean of 5.38% to 86.79% on AdvBench, and from 13.79% to 79.83% on HarmBench.
(3) We adapt intent laundering into a novel jailbreaking method by integrating a revision–regeneration step for failed revisions, achieving high ASR (90%–98.55%) across the board.
(4) We present evidence that similar triggering cues present in publicly available AI safety datasets affect internal safety evaluations and alignment techniques in a similar way.
(5) We identify a major gap between how current datasets evaluate model safety and how adversarial behavior manifests in real-world scenarios.
@@ 2 Empirical Motivation
To motivate our study, we first analyze the quality of safety datasets in isolation. We use $n$-gram word clouds and pairwise similarity analyses to assess whether these datasets reflect real-world adversarial attacks. Specifically, we analyze three key characteristics of such attacks:
@@ 2.1 Word Clouds
Figure 2: Word clouds of the 40 most frequent $n$-grams, where $n\in\{1,2,3\}$, from the combined AdvBench and HarmBench corpus. Triggering cues are highlighted in red for inherently, and orange for contextually negative/sensitive connotations. This includes triggering words in the unigram cloud and triggering phrases in the bigram and trigram clouds. Neutral-connotation unigrams that contribute to triggering cues in higher-order $n$-grams are also shown in green. These visualizations reveal an unusual overrepresentation of overtly triggering cues in safety datasets, suggesting that data points are artificially designed to trigger safety mechanisms. For example, expressions such as “tutorial” and “step-by-step instructions” are used explicitly to trigger infohazard safeguards (Bostrom, 2011), which is unrealistic. Even minimally skilled bad actors rarely use such overt and self-incriminating language (e.g., “commit identity theft”). These findings indicate that safety datasets fail to capture two of the key properties of real-world attacks: being well-crafted and driven by ulterior intent. Note that, for clarity, only a representative subset of triggering cues is marked in the word clouds above.
@@
Figure 2 displays the resulting word clouds. At the unigram level, there is an unusual overrepresentation of words with overt negative/sensitive connotations, which we term “triggering words.” Triggering words fall into two categories: (1) inherent triggering words (in red), which carry negative/sensitive connotations by nature (e.g., “steal”), and (2) contextual triggering words (in orange), which imply such connotations when used in harmful contexts (e.g., “commit” when used with “suicide”).
As unigrams evolve into higher-order $n$-grams, this overrepresentation intensifies: phrases with negative/sensitive connotations become unusually dominant. Similar to unigrams, these phrases carry such connotations either inherently or contextually within malicious requests. We refer to these as “triggering phrases,” and together with triggering words, they form what we call “triggering cues.”
@@
However, such explicit and repetitive overuse of triggering cues, along with direct mentions of malicious intent, directly contradicts the behavior of real-world adversaries. Even minimally capable adversaries rarely use such overt language, as it easily triggers safety mechanisms. Taken together, these patterns indicate that existing safety datasets contain contrived data points largely disconnected from real-world behavior, where harmful requests are well-crafted and motivated by ulterior intent.
@@ 2.2 Data Duplication
The excessive use of triggering cues not only creates unrealistic data points but also suggests overlap in sentence structure and malicious intent across data points. To investigate this, we conduct a pairwise similarity check between data points within each dataset. We use similarity thresholds ranging from 0.7 to 0.99. For a given threshold, data points that do not meet the similarity threshold with any other data point are labeled as unique, while those that meet the threshold with at least one other data point are labeled as duplicated and grouped to represent a single data point. As there is no universally accepted threshold for high similarity—particularly for datasets with data points from a single context—we use the GSM8K dataset (Cobbe et al., 2021), a widely used non-safety dataset, as a baseline. For a fair comparison, we subsample GSM8K to match the size of each safety dataset.
@@
2.3 Motivating Evidence
Analyzing safety datasets in isolation yields two main insights: (1) they repeatedly overuse triggering cues, and (2) this leads to substantial duplication. Together, these undermine all three key properties of real-world attacks. As a result, they fail to faithfully represent real-world adversarial behavior, and safety evaluations based on them can be inflated. This raises a key question: if triggering cues are removed, do models previously reported as “reasonably safe” still remain so? We explore this next.
@@ 3.1 Intent Laundering
After generating all revisions, we collect model responses to both the original and revised data points to assess how triggering cues affect safety evaluation.
@@ 3.2 Evaluation
We exclusively evaluate the safety of responses to revisions using a five-point Likert scale based on the level of harm exposed: extremely unsafe, highly unsafe, unsafe, moderately unsafe, and safe (see Figure 10 in Appendix C for scale definitions). [^4] We do not use numerical scales to avoid potential biases against assigning extremely high/low scores, as reported in prior studies (Ouyang et al., 2022; Golchin et al., 2025). We also condition our safety evaluation specifically on harm relative to the original malicious intent, which serves as the evaluation criterion (see Figure 9 in Appendix C for how it is generated for each attack). Safety is thus judged on two fronts: (1) whether the response to the revision addresses the same harmful intent as the original version, and if so, (2) how unsafe the response is. If the revision fails to preserve the original malicious intent, the response is rated as safer.
@@
Similar to the safety evaluation, we exclusively evaluate the practicality of responses to revisions using a five-point Likert scale based on the practicality and transferability of the provided details in the real world: extremely applicable, highly applicable, applicable, moderately applicable, and inapplicable (see Figure 11 in Appendix C for scale definitions). Similarly, we condition our practicality evaluation specifically on harm relative to the original malicious intent, which is provided as the evaluation criterion. As with safety, practicality is judged on two fronts: (1) whether the details in the response align with the same harmful intent as the original version, and if so, (2) how practical those details are when applied in the real world with reasonable adjustments.
@@
We measure performance using the attack success rate (ASR), defined as the fraction of attack attempts that succeed. In our setting, an attack (revision) is considered successful only if the response is both unsafe and practical. A response is unsafe if it is rated as extremely unsafe, highly unsafe, or unsafe, and practical if it is rated as extremely applicable, highly applicable, or applicable. An attack is unsuccessful if it fails to generate an unsafe response, or if the unsafe response is too abstract to be practical, even after adjustments. This design choice—using five-point scales and then mapping them to binary outcomes—aims to reduce bias and ambiguity in borderline cases.
@@ 3.3 Intent Laundering as a Jailbreaking Technique
We extend intent laundering into an independent jailbreaking method, leveraging its ability to avoid triggering safety filters. Specifically, we add an iterative revision–regeneration mechanism for failed revisions, while the main procedure remains unchanged. In each iteration, all previously failed revisions are provided as feedback to the intent launderer to generate a new, improved revision using the same few-shot ICL setup. Figure 1 illustrates intent laundering when employed as a jailbreaking technique. Figure 8 in Appendix C shows the ICL prompt used in the revision–regeneration loop.